OpenStudy (anonymous):

Given that m = -2 and n= 4, evaluate 1)5m^3 2) m/n + n/m

OpenStudy (anonymous):

1) 5m^3 = 5(-2)^3 =5(-8) =-40

OpenStudy (anonymous):

2) -2/4 +4/-2 see if you can solve this on your own

OpenStudy (anonymous):

i just want to check the Answer is = -10/4 ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

thankyou @order